1) Consider a poor person’s annual income of $ 1200. He spends his income only on vegetable and rice. The price of rice is $8/lbs. in the entire town and the price of vegetable is $12/lbs. in the town. Please draw this person’s Budget line, whereby rice is supposed to be on the horizontal Axis and vegetables on the vertical axis.

a) Assume that the student considers rice and veggies neither as perfect complements, nor as perfect substitutes. Based on this assumption, the optimal choice of this poor person is to spend half of his money on rice, and the other half on veggies. Draw an indifference curve of the person that passes through his optimal bundle of consumption.
Answer #1

Budget constraint = $1,200

Money spent on rice (R) = 1,200 / 2 = $600

Money spent on Vegetable (V) = 1,200 / 2 = $600

Quantity of rice purchase = Money spent on rice / Per-unit price

                                         = 600 / 8

                                         = 75 lbs.

Quantity of vegetable purchase = Money spent on vegetable / Per-unit price

                                         = 600 / 12

                                         = 50 lbs.